# Piezoelectric actuation for integrated photonics

### v1

@@ 6.1 Microwave to Optical Converter
Superconducting (SC) circuits and qubits have experienced a rapid progress in recent decades, featuring a full set of microwave quantum gates with high fidelity [328]. Although the computation supremacy over the classical computer has been demonstrated [329], the requirement of working at cryogenic temperature limits its scalability, due to the finite cooling power of current dilution refrigerators and the limited space imposed by the demand coaxial cables for the driving and readout of SC qubits. These challenges prevent the realization of fault-tolerant quantum computing in a foreseeable future, which requires millions of qubits [330]. A potential solution is to interconnect multiple quantum processors hosted in separate dilution refrigerators. Although the connection via coaxial cables has been demonstrated [331], they must be kept at cryogenic temperatures to maintain a low thermal noise, which is expensive and not scalable. However, optical photons possess frequencies four orders of magnitude higher than the microwave, leading to negligible thermal noise at room temperature. Indeed, the realization of the optical quantum key distribution (QKD) [332] has demonstrated its capability to maintain quantum superposition and entanglement at room temperature. Moreover, optical fiber telecommunication systems have become one of the major infrastructures which can be readily upgraded for building large-scale quantum processor and quantum network. Therefore, conversion between microwave and optical photons with high quantum efficiency has become an active research area recently, and promising progress has been made.
@@
For an ideal microwave to optical converter in quantum transduction, it should have high quantum efficiency with low added noise and wide transduction bandwidth. Although classical optical modulator based on electro-optic effect has been well studied, due to the large frequency and energy difference between microwave and optical photons, it is nontrivial to convert directly with high quantum efficiency through the weak nonlinearity of most optical materials. Nevertheless, electro-optic converter has been explored in various platforms, such as integrated AlN [187, 333] and LiNbO3 [167, 311, 334] photonic circuits, as well as bulk LiNbO3 disk [335, 336, 337]. Quantum transduction via an intermediary mode has attracted much attention, such as mechanical mode [338, 339], magnon mode [340, 341], and atomic transition [342]. Among them, the mechanical mode has been more popular due to its high efficiency, ease of fabrication, and compatibility with SC qubits.
@@
There are mainly two ways of coupling mechanics with electrical signals, including electrostatic and piezoelectric. The highest conversion efficiency has been achieved in electrostatically coupled mechanical membrane, from 10 [338] to 47 [343], reaching the 50 efficiency required for having finite quantum capacity in quantum communication [344]. However, since the membrane operates at MHz frequencies, it only supports a limited bandwidth, and the added noise is high even at cryogenic temperature. Progresses have been made recently in electrostatically exciting Si optomechanical crystal (OMC) which supports GHz mechanical modes [345, 309]. Despite of being applicable to most materials, the electrostatic coupling is still weaker than the piezoelectric force. In that sense, piezo-optomechanical transducer has drawn much attention in coherent conversion between microwave and optics. The variety of mechanical modes as presented in the previous chapter for AOM have also been explored in quantum transducer, from SAW [15, 306], BAW [346, 347, 348], to OMC [323, 284, 285].
@@
where with the intra-cavity pump photon number. The first and second terms in the above equation correspond to the beam-splitter and two-mode squeezing interaction, respectively. Depending on the detuning of the pump laser (assuming resolved sideband ), one interaction will dominate the other, with red detuning for beam-splitter and blue detuning for two-mode squeezing [186]. The quantum efficiency is defined as the ratio of the output optical photon to input microwave photon in up-conversion, and vice versa in down-conversion. Practically, the total efficiency consists of the coupling efficiency from the off-chip (microwave or optical) sources / detectors to the on-chip waveguides and the on-chip conversion efficiency . The can serve as a figure of merit for comparing different works and can be derived from the Heisenberg-Langevin equations of motion (see Eqs. 20-21) as [348]:
@@
The Hamiltonian that describes the interaction between the microwave and mechanical resonators is . The conversion efficiency by including the microwave resonator can be obtained similarly from the equations of motion as [309, 284]:
@@
Since the conversion efficiency is proportional to the intracavity pump photon number, and thus the on-chip optical pump power, high efficiency can be achieved by increasing the pump power. However, the coherent quantum transduction requires the added noise to be smaller than one. Therefore, the maximum power is ultimately limited by the added noise resulting from optical absorption, which has two main negative effects on the transducer. One is the generation of quasi-particles on the superconductor, and the other is the increased thermal occupancy of the mechanical mode. The quasi-particle is generated from broken copper pairs which can come either directly from the light absorption or through thermal heating [349, 285]. It has been experimentally examined that the quasi-particle could lower the lifetime of superconducting qubits [285] and shift the frequency and increase the intrinsic loss of the superconducting resonator [349]. The quasi-particle can be effectively mitigated by reducing the time of optical illumination through optical pulses [349]. Also, the use of superconductors with shorter quasi-particle lifetime could accelerate the recovery time of the superconductor, such as Nb and NbTiN which have quasi-particle lifetime on the order of nanoseconds [353].
@@
The dependence of and on for an 1D OMC is measured as shown in Fig. 44(c-d). It can be seen the optical absorption not only increase the thermal occupancy but also the dissipation of the mechanical resonator.
@@
For most of the optomechanical system in quantum transduction, it works in the resolved sideband regime where the mechanical frequency is much larger than the optical linewidth. This inevitably leads to low pump efficiency when the laser is red or blue detuned by where most of the pump light is reflected by the optical cavity. Moreover, it is shown that the heating due to the pump light in the coupling waveguide is comparable to that in the cavity [352]. To circumvent this problem, it has been proposed to include a second optical mode that aligns with the pump frequency, as shown in Fig. 42. In this way, the same intra-cavity photon number can be achieved with much smaller pump power. This approach is initially proposed and demonstrated in electro-optic converters [359, 360, 167, 311, 334, 333]. Recently, it has been implemented in HBAR [348] and OMC [361] schemes, where a second optical cavity is evanescently coupled to the cavity with optomechanical interaction. The two cavities form hybrid symmetric and asymmetric modes with spacing equal to the mechanical frequency. Depending on the application, either the lower or higher frequency mode can be pumped to drive the quantum transduction.
Scheme | SAW | BAW | OMC | OMC | OMC | OMC | OMC | EO | EM |
@@
Year | 2019 | 2023 | 2020 | 2022 | 2023 | 2023 | 2023 | 2020 | 2023 |
Material platform | LiNbO3 | AlN on Si3N4 | LiNbO3 | GaP | AlN on Si | LiNbO3 on Si | LiNbO3 on Si | LiNbO3 | Si |
(GHz) | 2 | 3.48 | 1.85 | 2.8 | 5 | 3.596 | 5.043 | 6.8 | 4.93 |
@@
Pump mode | CW | CW & Pulse | CW | CW & Pulse | Pulse | CW & Pulse | CW & Pulse | CW | CW |
Optical power | 1 mW | 126 mW | 3.3 W | 0.5 W | 6.2 | 10.2 W | 1 W | 1.2 mW | - |
1.04 | 0.3 | 0.13 | 0.268 | ||||||
@@
The comparison of microwave to optical converters realized in different schemes is summarized in Table 7. It can be seen the OMC has an optomechanical coupling strength orders of magnitude larger than SAW and BAW due to the tight mode confinement, which leads to large cooperativity under modest optical power. However, the sub-micron mechanical mode makes it hard to be driven directly through microwave, showing a small . The development of high-impedance microwave resonator in recent works [316, 317, 349] helps boost the drive efficiency. On the other hand, the advantages of SAW and BAW include lower optical losses and better optical power handling capability. Nevertheless, OMC is the most promising platform so far that would meet most of the requirements for quantum transduction. The highest efficiency of 5 has been achieved on the platform of LiNbO3 on Si OMC with only 10 W optical pump power.
@@
Although the quantum transduction efficiency of current converters is still far below the requirement for building the quantum internet [362, 344], the superconducting quantum processor can already benefit from the development of this technology in the near furture. One potential application is the classical optical control and readout of superconducting (SC) qubit. Conventionally, each qubit is connected via the coaxial cable to the control electronics at room temperature. However, this poses challenges on the scaling of the number of qubits which is limited to a thousand qubits based on current technology [363]. Apart from the space constraint, the heat load from coaxial cables would ultimately overwhelm the cooling power of the dilution refrigerator [364]. Optical fibers made from silica, on the other hand, show negligible thermal conductivity at cryogenic temperatures. Also, the ultra-wide optical bandwidth allows signal multiplexing in a single fiber. In that sense, replacing the coaxial cable with the photonic link could become an attracting solution to large-scale quantum computer in the future. As shown in Fig. 45(a), the converter down-converts the input light into microwave signals to drive and readout the qubit, and the readout signal is then up-converted into light for qubit measurement. It has recently been demonstrated that the optically generated microwave via photodiodes can faithfully serve as both drive and readout signal for an SC qubit [363]. Another work realized the optical readout of SC devices using off-the-shelf electro-optic modulators [365]. Based on these proofs of the concept, the conversion efficiency and added noise of the photonic link can be further improved using the-state-of-the-art microwave to optical converters. While most recent works on this topic have been realized on electro-optical converters [366, 367, 368], the optical multi-shot readout of SC qubit has been demonstrated on piezo-optomechanical transducer with high fidelity [369]. In the future, the investigation of heat load from these converters and the demonstration of optical multiplexing could enhance the scalability of this technology as a photonic link for SC circuits.
@@
Aside from the classical control and readout of qubits, it has been proposed that current converter technology can be readily applied in the optically heralded remote entanglement between distant SC systems [370] and the qubit states can be transmitted through the quantum teleportation [371], which has less strict requirement on the conversion efficiency. Following the well-known DLCZ protocol [372], the diagram of the proposed architecture is illustrated in Fig. 45(b). In stead of red detuned optical pump, the light is pumped at the blue side, which generates an entangled microwave-optical photon pair through an optomechanical parametric down conversion process [373, 374]. Interestingly, the transducer preferably works in low-cooperativity regime such that the probability of multi-photon events is kept low [368]. In this way, by passing the optical photons generated from separate SC systems through a beamsplitter to erase the path information, the detection of a single optical photon heralds the entanglement between microwave photons located in each SC circuits. The entangled microwave states can then be transferred to corresponding SC qubits in each node to build the quantum network. As a prerequisite, the generation of correlated microwave-optical photon pair and optically heralded non-classical microwave photon have been demonstrated in piezo-optomechanical transducers [317, 316]. Further improvement of the converter’s performance and reducing the losses in the system could increase the entanglement generation rate and fidelity, paving the way to the quantum network in a foreseeable future.
@@ 6.2 Magnetic-Free Optical Isolator
Apart from the SAW AOM, HBAR has also been utilized in demonstrating an optical isolator on the Si3N4 photonic circuit [228], as shown in Fig. 46(e). Differently from the SAW, HBAR excites the out-of-plane acoustic waves, which carry momentum perpendicular to the optical waveguide. To generate the in-plane momentum, it is proposed to include three discrete HBAR modulators that are driven by relative phases rotating in a circle. In this way, an effective rotational wave can be synthesized along the optical microring resonator. The resonator supports a pair of TE and TM modes with orthogonal polarization, which are then coupled through the modulation. Despite of the entirely different material systems and the choice of optical and mechanical modes, similar optical isolation performance is achieved with 10 dB isolation and 0.8 dB insertion loss, which demonstrates the universality of the dynamic modulation approach. Furthermore, the actuator phases can be reconfigured to switch the rotation direction such that the isolation direction can be controlled in situ, as illustrated in Fig. 46(f). The modulation sideband in the forward direction can be easily separated by a polarizing beamsplitter. The isolation of optical pulses is also demonstrated in Fig. 46(g). Despite of the compact size, the isolation bandwidth is limited by the optical linewidth to hundreds of megahertz to gigahertz. Although bandwidth can be increased by adding extra optical losses, the required mode splitting also increases, which leads to larger RF power. Nevertheless, the narrow bandwidth is sufficient in protecting laser systems and suitable for rejecting back-reflection noise for microwave-to-optical quantum transducers.
@@
Table 8 summarizes and compares the performances of optical isolator realized on different platforms. It can be seen that the scheme of spatio-temporal modulation has become one of the most promising methods in terms of compatibility with current photonic circuits, and the development of integrated AOM technology facilitates the recent demonstration of optical isolator on Si3N4 [228], Si [300], and LiNbO3 [140] photonics. While the ring-based structure is more compact, the waveguide structure supports wider bandwidth, which can be chosen according to specific application. Thus far, the widest bandwidth is achieved on the SAW modulated Si waveguides [300], and the highest isolation is demonstrated on the LiNbO3 electro-optic phase modulator [391]. The required RF power is on the order of 100 mW. These performances can be further improved by engineering the dispersion of optical waveguides, studying different pair of modes (e.g. TE and TM [403]), and exploiting new piezoelectric material.

### v2

@@ 6.1 Microwave to Optical Converter
Superconducting circuits and qubits have experienced a rapid progress in recent decades, featuring a full set of microwave quantum gates with high fidelity [344]. Although the prototype of quantum processing unit (QPU) has been demonstrated [345], the requirement of working at cryogenic temperature limits its scalability, due to the finite cooling power of current dilution refrigerators and the limited space imposed by the need for coaxial cables to drive and readout of SC qubits. These challenges prevent the realization of fault-tolerant quantum computing in a foreseeable future, which requires millions of qubits [346]. A potential solution is to interconnect multiple quantum processors hosted in separate dilution refrigerators. Although the proof-of-concept experiment has demonstrated the connection via coaxial cables [347], they must be kept at cryogenic temperatures to maintain a low thermal noise, which is expensive and not scalable. However, optical photons possess frequencies four orders of magnitude higher than the microwave, leading to negligible thermal noise at room temperature. Indeed, the realization of the optical quantum key distribution (QKD) [348] has demonstrated its capability to maintain quantum superposition and entanglement at room temperature. Moreover, optical fiber telecommunication systems have become one of the major infrastructures which can be readily upgraded for building large-scale quantum processor and quantum network. Therefore, conversion between microwave and optical photons with high quantum efficiency has become an active research area recently, and promising progress has been made.
@@
For an ideal microwave to optical converter in quantum transduction, it should have high quantum efficiency with low added noise and wide transduction bandwidth. Although classical optical modulator based on electro-optic effect has been well studied, due to the large frequency and energy difference between microwave and optical photons, it is nontrivial to convert directly with high quantum efficiency through the weak nonlinearity of most optical materials. Nevertheless, electro-optic conversion has been explored in various platforms, such as integrated AlN [199, 349] and LiNbO3 [179, 327, 350] photonic circuits, as well as bulk LiNbO3 disk [351, 352, 353]. Quantum transduction via an intermediary mode, such as mechanical mode [354, 355], magnon mode [356, 357], and atomic transition [358] has attracted much attention. Among them, the mechanical mode has been more popular due to its high efficiency, ease of fabrication, and compatibility with SC qubits.
@@
There are mainly three ways of coupling mechanics with microwave signals, including radiation pressure, electrostatic and piezoelectric. The highest conversion efficiency has been achieved in electromechanically coupled mechanical membrane via radiation pressure, from 10 [354] to 47 [359], approaching the 50 efficiency required for having finite quantum capacity in quantum communication [360]. However, since the membrane operates at MHz frequencies, it only supports a limited bandwidth, and the added noise is high even at cryogenic temperature. Progresses have been made recently in electrostatically exciting Si OMC which supports GHz mechanical modes [325]. Despite being applicable to most materials, the electrostatic coupling is still weaker than the piezoelectric force. In that sense, piezo-optomechanical transducer has drawn much attention in coherent conversion between microwave and optics. The variety of mechanical modes as presented in the previous chapter for AOM have also been explored in quantum transducer, from SAW [15, 322], BAW [361, 362, 363], to OMC [339, 297, 298].
@@
where with the intra-cavity pump photon number. The first and second terms in the above equation correspond to the beam-splitter and two-mode squeezing interaction, respectively. Depending on the detuning of the pump laser (assuming resolved sideband ), one interaction will dominate the other, with red detuning for beam-splitter and blue detuning for two-mode squeezing [198].
The microwave-to-optical conversion relies on the beam-splitter interaction where the annihilation of an input microwave photon leads to the generation of an output optical photon, and vice versa. The quantum efficiency is defined as the ratio of the output optical photon flux to input microwave photon flux in up-conversion, and vice versa in down-conversion. Practically, the total efficiency consists of the coupling efficiency from the off-chip (microwave or optical) sources / detectors to the on-chip waveguides and the on-chip conversion efficiency . The can serve as a figure of merit for comparing different works and can be derived from the Heisenberg-Langevin equations of motion (see Eqs. 21-22) as [363]:
@@
The Hamiltonian that describes the interaction between the microwave and mechanical resonators is . The conversion efficiency by including the microwave resonator can be obtained similarly from the equations of motion as [325, 297, 366]:
@@
Since the conversion efficiency is proportional to the intracavity pump photon number, and thus the on-chip optical pump power, high efficiency can be achieved by increasing the pump power. However, the coherent quantum transduction requires the added noise to be smaller than one. Therefore, the maximum power is ultimately limited by the added noise resulting from optical absorption, which has two main negative effects on the transducer. One is the generation of quasi-particles on the superconductor, and the other is the increased thermal occupancy of the mechanical mode. The quasi-particle is generated from broken Cooper pairs which can come either directly from the light absorption or through thermal heating [364, 298]. It has been experimentally examined that the quasi-particle could lower the lifetime of SC qubits [298] and shift the frequency and increase the intrinsic loss of the superconducting resonator [364]. The quasi-particle can be effectively mitigated by reducing the time of optical illumination through optical pulses [364]. Also, the use of superconductors with shorter quasi-particle lifetime could accelerate the recovery time of the superconductor, such as Nb and NbTiN which have quasi-particle lifetime on the order of nanoseconds [369].
@@
The dependence of and on for an 1D OMC is measured as shown in Fig. 44(c-d). It can be seen that the optical absorption not only increases the thermal occupancy but also the dissipation of the mechanical resonator.
@@
For most of the optomechanical system in quantum transduction, it works in the resolved sideband regime where the mechanical frequency is much larger than the optical linewidth. This inevitably leads to low pump efficiency when the laser is red or blue detuned by where most of the pump light is reflected by the optical cavity. Moreover, research indicates that the reflected light can constructively interfere with the incoming pump light, resulting in light absorption and heating within the coupling waveguide comparable to that in the cavity [368]. In this case, in the phenomenological model in Eq. 36 needs to be replaced by in order to include the heating from coupling waveguide, where is the input optical power [375]. In the demonstration in Ref. [368], the coupling waveguide is butt coupled with the OMC such that they are mechanically connected and the heating from the waveguide can be directly conducted to the OMC. The prefactor is experimentally measured to be 15 W-1 which largely increases the thermal occupation [368]. To circumvent this problem, it has been proposed to mechanically detach the waveguide from the OMC though side coupling [375]. equals to zero has been demonstrated with good theoretical fit with the experimental results [375].
A second method to mitigate the waveguide heating is to include a second optical mode that aligns with the pump frequency, as shown in Fig. 42. In this way, the same intra-cavity photon number can be achieved with much smaller input optical power , which largely suppresses the heating contribution from the waveguide. This approach is initially proposed and demonstrated in electro-optic converters [376, 377, 179, 327, 350, 349]. Recently, it has been implemented in HBAR [363] and OMC [378] schemes, where a second optical cavity is evanescently coupled to the cavity with optomechanical interaction. The two cavities form hybrid symmetric and asymmetric modes with spacing equal to the mechanical frequency. Depending on the application, either the lower or higher frequency mode can be pumped to drive the quantum transduction.
Scheme | SAW | BAW | OMC | EO | EM | ||||
@@
Year | 2019 | 2023 | 2020 | 2022 | 2023 | 2023 | 2023 | 2021 | 2023 |
Piezo Material | LiNbO3 | AlN | LiNbO3 | GaP | AlN | LiNbO3 | LiNbO3 | - | - |
Optical Material | LiNbO3 | Si3N4 | LiNbO3 | GaP | Si | Si | Si | LiNbO3 | Si |
(GHz) | 2 | 3.48 | 1.85 | 2.8 | 5 | 3.596 | 5.043 | 7.836 | 4.93 |
@@
Pump mode | CW | CW & Pulse | CW | CW & Pulse | Pulse | CW & Pulse | CW & Pulse | CW & Pulse | CW |
Optical power | 1 mW | 126 mW | 3.3 W | 0.5 W | 6.2 | 10.2 W | 1 W | 20 mW | - |
1.04 | 0.3 | 0.13 | 0.041 | 0.268 | |||||
@@
The comparison of microwave to optical converters realized in different schemes is summarized in Table 7. It can be seen that the OMC has an optomechanical coupling strength orders of magnitude larger than SAW and BAW due to the tight mode confinement, which leads to large cooperativity under modest optical power. However, the sub-micron mechanical mode makes it hard to be driven directly through a large microwave feed line, resulting in a small . The development of high-impedance microwave resonators in recent works [332, 333, 364] helps boost the drive efficiency. On the other hand, the advantages of SAW and BAW include lower optical losses and better optical power handling capability. Also, the electromechanical extraction efficiency and phonon injection efficiency for SAW and BAW are much higher. Even without additional microwave resonators, critically coupled HBAR has been demonstrated by engineering the acoustic impedance [379]. Nevertheless, OMC is the most promising platform so far that would meet most of the requirements for quantum transduction. The highest efficiency of 5 has been achieved with LiNbO3 on Si OMC with only 10 W optical pump power [333]. The added noise referred to the input is measured to be for optical-to-microwave conversion and 181 for microwave-to-optical conversion [333].
@@
Although the quantum transduction efficiency of current converters is still far below the requirement for building the quantum internet [380, 360], the superconducting quantum processor can already benefit from the development of this technology in the near furture. One potential application is the classical optical control and readout of SC qubits. Conventionally, each qubit is connected via a coaxial cable to the control electronics at room temperature. However, this poses challenges on the scaling of the number of qubits which is limited to a thousand qubits based on current technology [381]. Apart from the space constraint, the heat load from coaxial cables would ultimately overwhelm the cooling power of the dilution refrigerator [382]. Optical fibers made from silica, on the other hand, show negligible thermal conductivity at cryogenic temperatures. Also, the ultra-wide optical bandwidth allows signal multiplexing in a single fiber. In that sense, replacing the coaxial cable with the photonic link could become an attracting solution to large-scale quantum computer in the future. As shown in Fig. 45(a), the converter down-converts the input light into microwave signals to drive and readout the qubit, and the readout signal is then up-converted into light for qubit measurement. It has recently been demonstrated that the optically generated microwave via photodiodes can faithfully serve as both drive and readout signal for an SC qubit [381]. Another work realized the optical readout of SC devices using off-the-shelf electro-optic modulators [383]. Based on these proofs of the concept, the conversion efficiency and added noise of the photonic link can be further improved using the-state-of-the-art microwave to optical converters. While most recent works on this topic have been realized on electro-optical converters [384, 385, 386], the optical multi-shot readout of SC qubit has been demonstrated on piezo-optomechanical transducer with high fidelity [387]. In the future, the investigation of heat load from these converters and the demonstration of optical multiplexing could enhance the scalability of this technology as a photonic link for SC circuits.
@@
Aside from the classical control and readout of qubits, it has been proposed that current converter technology can be readily applied for the optically heralded remote entanglement between distant SC systems [388] and the quantum teleportation based transduction between microwave and optical photons [389], which has less strict requirement on the conversion efficiency. Following the well-known DLCZ protocol [390], the diagram of the proposed architecture is illustrated in Fig. 45(b). In stead of red detuned optical pump, the light is pumped at the blue side, which generates an entangled microwave-optical photon pair through an optomechanical parametric down conversion process [391, 392]. Interestingly, the transducer preferably works in a low optomechanical cooperativity regime such that the probability of multi-photon events is kept low [386]. In this way, by passing the optical photons generated from separate SC systems through a beamsplitter to erase the path information, the detection of single optical photons in both outputs of the beam splitter heralds the entanglement between microwave photons located in each SC circuits. The entangled microwave states can then be transferred to corresponding SC qubits in each node to build the quantum network. Different entanglement schemes have been proposed and the entanglement fidelity has been theoretically analyzed based on practical experimental parameters, such as time-bin [391] and frequency-bin [392] encoded entanglement. To have a high entanglement fidelity, the system prefers large electromechanical cooperativity and microwave extraction efficiency to suppress the microwave photon loss during the transduction [391, 392, 389]. Large is relatively easier to reach than since the piezo-electromechanical coupling rate is generally on the similar order of magnitude with the microwave and mechanical resonators’ linewidths (see Table 7). On the other hand, entanglement purification protocol and quantum error correction codes can be utilized to improve the entanglement fidelity at the presence of microwave loss [388, 393]. Ref. [391] proposes to improve the fidelity by postselecting the quantum states via a joint parity measurement of the SC qubits in a time-bin encoded entanglement. Preliminary experiments have demonstrated the generation of correlated microwave-optical photon pairs and optically heralded non-classical microwave photons in piezo-optomechanical transducers [333, 332]. Further improvement of the converter’s performance and reduction of the losses in the system could increase the entanglement generation rate and fidelity, paving the way to the quantum network in a foreseeable future.
@@ 6.2 Magnetic-Free Optical Isolator
Apart from the SAW AOM, HBAR has also been utilized in demonstrating an optical isolator on the Si3N4 photonic circuit [55], as shown in Fig. 46(e). Differently from the SAW, HBAR excites the out-of-plane acoustic waves, which carry momentum perpendicular to the optical waveguide. To generate the in-plane momentum, it is proposed to include three discrete HBAR modulators that are driven by relative phases rotating in a circle. In this way, an effective rotational wave can be synthesized along the optical microring resonator. The resonator supports a pair of TE00 and TM00 modes with orthogonal polarization, which are then coupled through the modulation. Despite of the entirely different material systems and the choice of optical and mechanical modes, similar optical isolation performance is achieved with 10 dB isolation and 0.8 dB insertion loss, which demonstrates the universality of the dynamic modulation approach. Furthermore, the actuator phases can be reconfigured to switch the rotation direction such that the isolation direction can be controlled in situ, as illustrated in Fig. 46(f). The modulation sideband in the forward direction can be easily separated by a polarizing beamsplitter. The isolation of optical pulses is also demonstrated in Fig. 46(g). Despite of the compact size, the isolation bandwidth is limited by the optical linewidth to hundreds of megahertz to gigahertz. Although bandwidth can be increased by adding extra optical losses, the required mode splitting also increases, which leads to larger RF power. Nevertheless, the narrow bandwidth is sufficient in protecting laser systems and suitable for rejecting back-reflection noise for microwave-to-optical quantum transducers.
@@
Table 8 summarizes and compares the performances of optical isolator realized on different platforms. It can be seen that the scheme of spatio-temporal modulation has become one of the most promising methods in terms of compatibility with current photonic circuits, and the development of integrated AOM technology facilitates the recent demonstration of optical isolator on Si3N4 [55], Si [316], and LiNbO3 [149] photonics. While the ring-based structure is more compact, the waveguide structure supports wider bandwidth, which can be chosen according to specific application. Thus far, the widest bandwidth is achieved on the SAW modulated Si waveguides [316], and the highest isolation is demonstrated on the LiNbO3 electro-optic phase modulator [410]. The required RF power is on the order of 100 mW. These performances can be further improved by engineering the dispersion of optical waveguides, studying different pair of modes (e.g. TE00 and TM00 [422]), and exploiting new piezoelectric material.